the sum of a two digit number and the number formed by reversing the order digits is 88. if difference of the digits is 2 and the units digit is greater. find the number
Answer by ankor@dixie-net.com(22740)

the sum of a two digit number and the number formed by reversing the order digits is 88.
if difference of the digits is 2 and the units digit is greater.
find the number
:
Let x = the 10's digit
Let y = the units
:
Write an equation for each statement:
:
"the sum of a two digit number and the number formed by reversing the order digits is 88."
(10x + y) + (10y + x) = 88
11x + 11y = 88
Simplify, divide by 11
x + y = 8
:
"if difference of the digits is 2 and the units digit is greater."
y - x = 2
we can use elimination with the 1st equation
y - x = 2
y + x = 8
--------------Adding eliminates x, find y
2y = 10
y = 5 is the units digit
then
5 - 2 = 3 is 10's digit
:
35 is the number
:
:
:
Check solution in the 1st statement:
the sum of a two digit number and the number formed by reversing the order digits is 88.
35 + 83 = 88